A camera flash consists of a 120 μF capacitor charged to 220V which discharges through a flashbulb with 5 Ω resistance. What is the time-constant τ of this RC circuit when the flash goes off? How much of the initial stored energy is discharged within 2τ? What is the equation for the voltage drop across the capacitor vs. time during the discharge? Make a graph showing this voltage vs. time.
Answer #1

Concept: use the discharging of capacitor through a resistance and the time constant of the discharging circuit as shown below

***************************************************************************************************
Check the answer and let me know immediately if you find something wrong or missing... I will rectify the mistakes asap if any
Answer #2

Solution:

  1. Time Constant (τ):

    τ=R×C=5Ω×120×106F=0.0006s=0.6ms

  2. Energy Discharged in 2τ:

    • Initial energy stored in the capacitor:

      E0=12CV2=12×120×106×(220)2=2.904J

    • Energy remaining after 2τ (since V=V0et/τ, at t=2τV=V0e2):

      E=12C(V0e2)2=E0e42.904×e40.053J

    • Energy discharged:

      Edischarged=E0E2.9040.053=2.851J(98.2% of E0)

  3. Voltage Drop Equation:

    V(t)=V0et/τ=220et/0.0006V

  4. Graph of Voltage vs. Time:

    • At t=τV=220e180.9V

    • At t=2τV=220e229.8V.

    • The voltage decays exponentially from 220V to near 0V over time.

    • Key points:


answered by: Harshwardhan kunal